OpenStudy (anonymous):

find maxima minima using 1st derivative test..... y=x^3+3x^2+3x y=(x-6)^2(9-x)

OpenStudy (anonymous):

=3x^2+6x+3 =3(x^2+2x+1) =(x+3)(x-1)=0 x=1 x=-3 is this correct?

OpenStudy (anonymous):

do you know what a derivative is?

OpenStudy (anonymous):

yes but i'm asking about maxima minima....

OpenStudy (anonymous):

right i understand, but to get maxima and minima according to the directions you provided we need to get the derivative

OpenStudy (anonymous):

the derivative of y=x^3+3x^2+3x is 3x^2+6x+3 right?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

then what should i do next?

OpenStudy (anonymous):

i'm confused in the part finding the critical points.

OpenStudy (anonymous):

well, my shadow will tell you

OpenStudy (anonymous):

let me help a little shessa. take the first derivative and set it = 0 find out where it is 0 and those are the critical point take the second derivative and set that =0

OpenStudy (anonymous):

the guy im shadowing will tell you what to do after you take the second derivative

OpenStudy (anonymous):

all you need is the first derivative though, since you are not looking concavity

OpenStudy (anonymous):

ok i dont mean to confuse. Critical points= first derivative shessa

OpenStudy (anonymous):

can you pls. answer the problems above so that i could clearly understand? pls..

OpenStudy (anonymous):

and now that i look at this function more closely , there are no maxima or minima

OpenStudy (anonymous):

can you pls. answer the problems above so that i could clearly understand? pls..

OpenStudy (anonymous):

how did you know?

OpenStudy (anonymous):

can you pls. answer the problems above so that i could clearly understand? pls..

OpenStudy (anonymous):

pls let me understand pls.

OpenStudy (anonymous):

x^3+3x^2+3x d/dx(x^3+3x^2+3x)=0 3x^2+6x+3=0 x^2+2x+1=0 (x+1)^2=0 x=-1

OpenStudy (anonymous):

this is saying first derivative= where the slope =0 because thats what we set it = to. so at x=-1 thats just where the slope is 0

OpenStudy (anonymous):

can you pls. answer the problems above so that i could clearly understand? pls..

OpenStudy (anonymous):

noxipro can you explain this part pls. x^2+2x+1=0 (x+1)^2=0 x=-1

OpenStudy (anonymous):

i understand in getting the derivative but how did it become x^2+2x+1=0 (x+1)^2=0 x=-1

OpenStudy (anonymous):

x^2+(a+b)x+ab=(x+a)(x+b) since 1=a and 1=b a+b=2 ab=1 so my a and b are both 1

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

am is that factoring?

OpenStudy (anonymous):

yes it is just factoring

OpenStudy (anonymous):

am what if its like this x^3-3x =3x^2-3 =3x^2-3=0 what is the critical points of this?

OpenStudy (anonymous):

3(x^2-1) is my factoring correct? then critical points are?

OpenStudy (anonymous):

jst divide everything by 3 because 0 is on the other side x^2-1=(x+1)(x-1)

OpenStudy (anonymous):

ah ok tnx.am the 2nd function can you help solve it?

OpenStudy (anonymous):

(x-6)^2(9-x) x^2-12x+36(9-x) 9x^2-12(9)x+36(9)-x^3+12x^2-36x d/dx=18x-12(9)-3x^2+24x-36 52x-144-3x^2=0 3x^2-42x+144=0 x^2-14x+48=0 (x-8)(x-6)=0 x=8 and x=6

OpenStudy (anonymous):

NoxiPro explain this part x^3+12x^2-36x pls..

OpenStudy (anonymous):

since you're foiling, you first multiply x^2 -12x +36 by 9, which is the first part of (9-x). thus: 9x^2-12(9)x+36(9) with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the second part of (9-x) is the -x. and you MUST carry the negative with it. you then multiply the original x^2-12x+36 by that -x. this gives you: x^2(-x)-12x(-x)+36(-x) which simplifies to: -x^3+12x^2-36x

OpenStudy (anonymous):

ah ok i got it now thank you very much